OpenStudy (anonymous):

Let v be the region in R3 consisting of all points (x,y,z) such that: 0≤x, 0≤y, 0≤z, x^2 + y^2 ≤ 1, x^2 + z^2 ≤ 1, y^2 + z^2 ≤ 1 Show that the volume of V equals: 2 (integral Phi from 0 to Pi/4)(integral r from 0 to 1) of ((1-r^2cos^2(Phi))^.5) r dr dPhi

OpenStudy (anonymous):

equation should look like this:\[2\int\limits_{0}^{Pi/4}\int\limits_{0}^{1}\sqrt{(1-r^2cos^2(\phi))}rdrd\phi\]

OpenStudy (anonymous):

@Zarkon , little help

OpenStudy (amistre64):

this looks like its the unit sphere located in the first octant

OpenStudy (anonymous):

its not the unit sphere

OpenStudy (amistre64):

\[\int_{0}^{pi/2}\int_{0}^{pi/2}\int_{0}^{1} dr\ d\theta \ d\phi\]

OpenStudy (amistre64):

x^2+y^2 = 1 is a circle in the xy plane x^2 + z^2 = 1 is a circle in the xz plane z^2 + y^2 = 1 is a circle in the zy plane

OpenStudy (amistre64):

imnot to confident in the spherical and cylindrical coordinates

OpenStudy (anonymous):

so you dont know how id go about showing that?

OpenStudy (amistre64):

i cent think of it right at the moment :) ill work at it and see what i can devise tho

OpenStudy (amistre64):

i still think its an eighth of the volume of a unit sphere 4pi/8 = pi/2 ... but ill see what i can come up with

OpenStudy (amistre64):

is your integrand given in terms of spherical or cylindrical ??

OpenStudy (anonymous):

unit sphere is given by x^2 + y^2 + z^2 = 1 , not what is given in the problem. intergrand is in spherical terms

OpenStudy (anonymous):

It is not the unit sphere. It is the intersection of the three cylinders described by the 3 inequalities in the first octant.

OpenStudy (anonymous):

they're not cylinders either, but parabolas opening up toward (0,0,0) on each axis

OpenStudy (anonymous):

They are cylinders. What is \[ x^2 + y^2 \le 1 \] It is a cylinder centered around the z-axis. Wait, I will graph it for you.

OpenStudy (amistre64):

pi/4 would split it in half and it is symmetrical in the region

OpenStudy (anonymous):

are you sure? if I set x = 1 then both y and z have to be equal to 0, which implies a parabola

OpenStudy (anonymous):

This is the later area of the cylinder \[ x^2 + y^2 \le 1 \] Its parametrization is p(r,z) = { cos t, sin t, z}}

OpenStudy (amistre64):

i would think that the term "phi" would indicate spherical and not cylindrical ...

OpenStudy (amistre64):

either way, i got no real idea how to "prove" it

OpenStudy (anonymous):

Suppose you are in the plane z=k, then \[x^2 + y^2 \le 1 \] it the graph of a disk of center the origin and of radius one in that plane. Take all the planes z=k, you get a cylinder

OpenStudy (anonymous):

It can't be a cylinder though, Suppose you set x = 1, by the given parameters, both y = 0 and z = 0. This is the vertex of a parabola, not a cylinder, since we cannot have x = 1 and y > 0 or z > 0

OpenStudy (anonymous):

See http://mathworld.wolfram.com/SteinmetzSolid.html For a solution to your problem. You have to adjust some parameters to do that.

OpenStudy (anonymous):

If x=1, then the point (1,0,0) is part of your solid. Why should it be a vertex of a parabola?

OpenStudy (anonymous):

because if x = 1, y has to be 0 and z has to be 0. y and z cannot take any other values.

OpenStudy (anonymous):

I agree, but that will give you one point (1,0,0) in your solid. What is wrong with that?

OpenStudy (anonymous):

The problem is that to make the disk at the top of the cylinder, we need the point (1,0,0) and all points (1,y,z) such that y^2 + z^2 = r^2 where r is the radius of the cylinder.

OpenStudy (anonymous):

Read (18)(19) and (20) of http://mathworld.wolfram.com/SteinmetzSolid.html take r=1 and divide by 8. There are 8 octants. You obtain the formula that you want.